OpenStudy (anonymous):

use lagrange multipliers to find the absolute extrema of f(x,y,z)= xyz subject to the constraint x^2+2y^2+4z^2=4.

OpenStudy (anonymous):

\[yz= 2x \lambda \] \[xz= 4y \lambda\] \[xy= 8z \lambda\] i always end up with incomplete critical points when solving problems like this..

OpenStudy (anonymous):

\[\begin{cases}yz=2x\lambda\\ xz=4y\lambda\\ xy=8z\lambda\\ x^2+2y^2+4z^2=4\end{cases}\] Multiply the first equation by \(x\), the second by \(y\), and the third by \(z\): \[\begin{cases}xyz=2x^2\lambda\\ xyz=4y^2\lambda\\ xyz=8z^2\lambda\\ x^2+2y^2+4z^2=4\end{cases}\] So now you have \[2x^2\lambda=4y^2\lambda=8z^2\lambda\\ x^2=2y^2=4z^2\] (assuming \(\lambda\not=0\)). Substituting into the fourth equation: \[x^2+x^2+x^2=4\\ x^2=\frac{4}{3}\\ x=\pm\sqrt\frac{4}{3}\] So you also have \[2y^2=\frac{4}{3}\\ y=\pm\sqrt\frac{2}{3}\] and \[4z^2=\frac{4}{3}\\ z=\pm\sqrt\frac{1}{3}\]